# Model Editing Harms General Abilities of Large Language Models: Regularization to the Rescue

### v3

@@ Abstract
Model editing is a technique that edits the large language models (LLMs) with updated knowledge to alleviate hallucinations without resource-intensive retraining. While current model editing methods can effectively modify a model’s behavior within a specific area of interest, they often overlook the potential unintended side effects on the general abilities of LLMs such as reasoning, natural language inference, and question answering. In this paper, we raise concerns that model editing’s improvements on factuality may come at the cost of a significant degradation of the model’s general abilities. We systematically analyze the side effects by evaluating four popular editing methods on three LLMs across eight representative tasks. Our extensive empirical experiments show that it is challenging for current editing methods to simultaneously improve factuality of LLMs and maintain their general abilities. Our analysis reveals that the side effects are caused by model editing altering the original model weights excessively, leading to overfitting to the edited facts. To mitigate this, a method named RECT (RElative Change in weighT) is proposed to regularize the edit update weights. Evaluation results show that RECT can significantly mitigate the side effects of editing while still maintaining over 94% editing performance.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Demonstration of model editing and its impact on the general abilities of LLMs. Although the factuality of the model has been improved, the general abilities of LLMs, such as question answering, sentiment analysis, dialogue, information extraction, have been substantially impaired after editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
As real-world knowledge is dynamically increasing and updating, existing large language models (LLMs) need to constantly incorporate the inherit knowledge and up-to-date information for life-long learning. Despite continual training, LLMs inevitably manifest hallucinations caused by missing, false or outdated knowledge embedded in their parameters (Zhang et al., 2023; Peng et al., 2023; Ji et al., 2023). Due to the intensive computational cost of retraining LLMs, researchers have increasingly focused on model editing (a.k.a., knowledge editing(Sinitsin et al., 2020; Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022b; Meng et al., 2022, 2023; Yao et al., 2023; Zhong et al., 2023; Ma et al., 2023; Zhang et al., 2024). This task is to efficiently modify a model’s behavior within a specific area of interest through targeted interventions without retraining.
At present, the assessment of editing methods typically involves evaluation along three critical dimensions (Yao et al., 2023). First, efficacy ensures the post-edited model can accurately recall the specific edited fact. Second, generalization assesses the model’s ability to recall the fact under diverse paraphrase prompts. Finally, locality checks if the model’s output for unrelated inputs remains consistent after editing. These multi-faceted criteria collectively contribute to a nuanced understanding of the effectiveness and robustness of editing methods.
@@
In light of the above issues, we systematically study if model editing hurts the general abilities of LLMs. This work studies model editing in the single- versus sequential-editing and instance- versus batch-editing settings. The edited models are evaluated on a variety of downstream tasks to see if there are any side effects on performance before and after editing. Extensive empirical experiments are conducted on four popular editing methods: KN (Dai et al., 2022), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023) applied to three representative LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). Eight representative tasks including reasoning (Cobbe et al., 2021), natural language inference (Dagan et al., 2005), open-domain QA (Kwiatkowski et al., 2019), closed-domain QA (Clark et al., 2019a), dialogue (Cui et al., 2020), summarization (Gliwa et al., 2019), named entity recognition (Sang & Meulder, 2003), and sentiment analysis (Socher et al., 2013) are employed to understand the impact of model editing on the general abilities of LLMs.
Experimental results show that existing LLMs are not robust to weight perturbations, and editing even a few parameters can significantly affect their general abilities. Strikingly, with a single pass of editing involving less than 1% parameters, LLaMA-1 (7B) exhibited a drastic performance degradation to nearly 0 on all the tasks we tried.
@@ 2 Related Work
Many studies have investigated model editing, including memory-based, meta-learning, and locate-then-edit (Wang et al., 2023; Yao et al., 2023). Memory-based methods do not modify model weights but store the editing facts with an external memory (Mitchell et al., 2022b; Zhong et al., 2023). Mitchell et al. (2022b) stored edits in a base model and learned to reason over them to adjust its predictions as needed. The latter two classes of methods are developed to directly modify the internal parameters of models, which is the focus of this paper. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update model parameters (Cao et al., 2021; Mitchell et al., 2022a). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022a) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly (Dai et al., 2022; Meng et al., 2022, 2023). In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module. Besides, some works investigate the evaluation paradigm for model editing (Zhong et al., 2023; Cohen et al., 2023; Ma et al., 2023; Li et al., 2023; Hase et al., 2023; Wu et al., 2023a; Gandikota et al., 2023; Wang et al., 2024). For example, Cohen et al. (2023) introduced the ripple effects of editing, suggesting that editing a particular fact implies that many other facts need to be updated. DeepEdit (Wang et al., 2024) is neuro-symbolic that decodes with constraints and can be flexibly applied to black-box LLMs. Additionally, recent works have also applied editing in various domains, such as changing model personality (Mao et al., 2023), editing multimodal models (Cheng et al., 2023), protecting users privacy (Wu et al., 2023b), etc.
Interpretability of Transformer
Many studies have attempted to explain and understand Transformer (Vaswani et al., 2017). Previous works focus on the function of self-attention layers and the differences between layers (Tenney et al., 2019; Vig & Belinkov, 2019; Voita et al., 2019; Clark et al., 2019b). For instance, Voita et al. (2019) found that the most confident attention heads played consistent linguistically-interpretable roles. More recently, a surge of works have investigated the role of feed-forward network (FFN) layers in Transformer, showing that FFN layers store factual and linguistic knowledge, which can be located in specific neurons and edited (Geva et al., 2021; Da et al., 2021; Meng et al., 2022; Dai et al., 2022). Besides, Geva et al. (2022) investigated the mechanism in which FFN layers update the inner representations of Transformer-based LMs. They propose that the FFN output can be viewed as a collection of updates that promote concrete concepts in the vocabulary space.
@@
A main difference between this work and previous studies should be highlighted. These approaches target at designing editing algorithms to improve or evaluation paradigms to assess the editing performance. In contrast, this study rethinks model editing and explores if current editing methods inadvertently cause the potential side effects on the underlying general abilities of LLMs. The contemporaneous work (Gupta et al., 2024) presents a similar finding that model editing at scale leads to catastrophic forgetting, but no mitigation method is proposed. To the best of our knowledge, this paper makes the first call for attention to side effects on a variety of tasks beyond editing performance by presenting a systematical evaluation of four editing methods on three LLMs covering eight tasks. Besides, we also analyze the causes of side effects, and propose a regularization method to prevent editing overfitting.
@@ 3 Preliminary
To evaluate the effectiveness of editing methods, previous works focus on evaluation along three dimensions (Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022, 2023).
Efficacy
Generalization
Locality
@@ 4.1 Evaluation Paradigm
Single-editing involves examining the efficacy and impact of making a single editing operation to a model. Specifically, it focuses on understanding how a model adapts to such a single alteration, and the implicit effect of such specific modifications on the overall performance. It is worth noting that a single editing operation can contain either only one editing instance or multiple ones in a batch, which is further discussed later. In practice, there are often situations where only a particular change is needed, so it’s crucial to understand how effectively the model integrates and preserves that individual edit. Therefore, evaluating the robustness to a single edit is crucial in determining its ability to retain the intended changes and overall performance.
In contrast to single-editing, multiple editing operations are conducted successively in sequential-editing (Huang et al., 2023). Similarly, each editing operation in sequential-editing can also contain either only one editing instance or multiple ones in a batch. Ideally, models should retain the changes from previous edits when carrying out a new one (Yao et al., 2023), which is decisive for the continual learning of future LLMs. Therefore, whether edited models can still maintain its general abilities after sequential editing is one of the important characteristics that should be considered. For this analysis, how the performance of edited models on a variety of tasks changes as the number of edits increases will be explored.
@@
The real world is ever-changing, so there is a huge amount of knowledge that needs to be dynamically added and updated into LLMs. Despite the effectiveness of many instance-editing methods (Dai et al., 2022; Meng et al., 2022; Ma et al., 2023), ultimately at most a few dozens of pieces of knowledge can be updated (Mitchell et al., 2022b), due to their relatively low but still non-negligible editing cost for a single instance. Since naive sequential applications of current state-of-the-art model editing methods fail to scale up (Meng et al., 2023), one may wish to update hundreds or thousands of facts simultaneously in batch-editing. Notably, batch-editing can also be coupled with both the single- or sequential-editing settings.
@@
Zero-shot learning aims to solve tasks without labeled training examples, and recent studies have demonstrated the superiority of LLMs for zero-shot learning (Brown et al., 2020; Wei et al., 2022; Chowdhery et al., 2023). Following these studies, we explore the zero-shot performance of unedited and edited models on a variety of tasks. Given a task instruction and a test problem that are concatenated as the input, the model is expected to generate a target text to address this problem. The instructions and input formats of different tasks are shown in Appendix A, which are taken from or inspired by Qin et al. (2023).
@@ 4.2 Evaluation Setup
Four popular editing methods were selected: (1) KN (Dai et al., 2022) involved identifying neurons linked to knowledge expression using gradient-based attributions and then enhancing the MLP layer by adding scaled embedding vectors to those specific neurons. (2) MEND (Mitchell et al., 2022a) learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form. (3) ROME (Meng et al., 2022) involved localizing factual knowledge in a specific Transformer MLP layer and then updating it by directly writing new key-value pairs into the MLP module. (4) MEMIT (Meng et al., 2023) expanded the capabilities of ROME by enabling the editing of large amounts of factual data through the updating of a sequence of MLP layers. It is notable that only MEND and MEMIT support batch-editing. All experiments were conducted using the EasyEdit tool (Wang et al., 2023), ensuring standardized and reproducible evaluation. All editing instances were randomly sampled from the editing dataset. Readers can refer to Appendix B for details of these editing methods.
@@
The popular model editing dataset Zero-Shot Relation Extraction (ZsRE) (Levy et al., 2017) used in previous work (Cao et al., 2021; Meng et al., 2022; Yao et al., 2023) was adopted in our experiments.
@@
Experiments were conducted on three LLMs: GPT-2 XL (1.5B) (Radford et al., 2019), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b).
@@
To extensively explore whether model editing has side effects on the general abilities of LLMs, eight representative tasks were adopted: (1) Reasoning on the GSM8K (Cobbe et al., 2021), and the results were measured by solve rate. (2) Natural language inference (NLI) on the RTE (Dagan et al., 2005), and the results were measured by accuracy of two-way classification. (3) Open-domain QA on the Natural Question (Kwiatkowski et al., 2019), and the results were measured by exact match (EM) with the reference answer after minor normalization as in Chen et al. (2017) and Lee et al. (2019). (4) Closed-domain QA on the BoolQ (Clark et al., 2019a), and the results were also measured by EM. (5) Dialogue on the MuTual (Cui et al., 2020), and the results were measured by selecting one best-matched response from four available candidates, denoted as Recall4@1 as in Lowe et al. (2015). (6) Summarization on the SAMSum (Gliwa et al., 2019), and the results were measured by the average of ROUGE-1, ROUGE-2 and ROUGE-L as in Lin (2004). (7) Named entity recognition (NER) on the CoNLL03 (Sang & Meulder, 2003), and the results were measured by entity-level F1-score. (8) Sentiment analysis on the SST2 (Socher et al., 2013), and the results were measured by accuracy of two-way classification.

### v4

@@ Abstract
Model editing is a technique that edits the large language models (LLMs) with updated knowledge to alleviate hallucinations without resource-intensive retraining. While current model editing methods can effectively modify a model’s behavior within a specific area of interest, they often overlook the potential unintended side effects on the general abilities of LLMs such as reasoning, natural language inference, and question answering. In this paper, we raise concerns that model editing’s improvements on factuality may come at the cost of a significant degradation of the model’s general abilities. We systematically analyze the side effects by evaluating four popular editing methods on three LLMs across eight representative tasks. Our extensive empirical experiments show that it is challenging for current editing methods to simultaneously improve factuality of LLMs and maintain their general abilities. Our analysis reveals that the side effects are caused by model editing altering the original model weights excessively, leading to overfitting to the edited facts. To mitigate this, a method named RECT is proposed to regularize the edit update weights by imposing constraints on their complexity based on the RElative Change in weighT. Evaluation results show that RECT can significantly mitigate the side effects of editing while still maintaining over 94% editing performance [^1] https://github.com/JasonForJoy/Model-Editing-Hurt.
Model Editing Harms General Abilities of Large Language Models: Regularization to the Rescue
Jia-Chen Gu1thanks:  Equal contribution., Hao-Xiang Xu211footnotemark: 1, Jun-Yu Ma2, Pan Lu1, Zhen-Hua Ling2, Kai-Wei Chang1, Nanyun Peng1 1University of California, Los Angeles 2University of Science and Technology of China {gujc,panbruin}@ucla.edu, {nh2001620,mjy1999}@mail.ustc.edu.cn zhling@ustc.edu.cn, {kwchang,violetpeng}@cs.ucla.edu
@@ 1 Introduction
Figure 1: Demonstration of model editing and its impact on the general abilities of LLMs. Although the factuality of the model has been improved, the general abilities of LLMs, such as question answering, dialogue, named entity recognition, sentiment analysis, have been substantially impaired after editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
As real-world knowledge is dynamically increasing and updating, existing large language models (LLMs) need to constantly incorporate the inherit knowledge and up-to-date information for life-long learning. Despite continual training, LLMs inevitably manifest hallucinations caused by missing, false or outdated knowledge embedded in their parameters Zhang et al. (2023); Peng et al. (2023); Ji et al. (2023). Due to the intensive computational cost of retraining LLMs, researchers have increasingly focused on model editing (a.k.a., knowledge editingSinitsin et al. (2020); Cao et al. (2021); Dai et al. (2022); Mitchell et al. (2022b); Meng et al. (2022, 2023); Yao et al. (2023); Zhong et al. (2023); Ma et al. (2023); Zhang et al. (2024). This task is to efficiently modify a model’s behavior within a specific area of interest through targeted interventions without resource-intensive model retraining.
At present, the assessment of editing methods typically involves evaluation along three critical dimensions Yao et al. (2023). First, reliability ensures the edited model can accurately recall the specific edited fact. Second, generalization validates the adaptability of the edited model by assessing the model’s ability to recall the fact under diverse paraphrase prompts. Finally, locality checks if the edited model’s output for unrelated inputs remains consistent after editing. These multi-faceted criteria collectively contribute to a nuanced understanding of the effectiveness and robustness of editing methods.
@@
In light of the above issues, we systematically study if model editing hurts the general abilities of LLMs. This work studies model editing in the single- versus sequential-editing and instance- versus batch-editing settings. The edited models are evaluated on a variety of downstream tasks to see if there are any side effects on performance before and after editing. Extensive empirical experiments are conducted on four popular editing methods: KN Dai et al. (2022), MEND Mitchell et al. (2022a), ROME Meng et al. (2022), and MEMIT Meng et al. (2023) applied to three representative LLMs: GPT-2 XL (1.5B) Radford et al. (2019), LLaMA-1 (7B) Touvron et al. (2023a), and LLaMA-2 (7B) Touvron et al. (2023b). Eight representative tasks including reasoning Cobbe et al. (2021), natural language inference Dagan et al. (2005), open-domain QA Kwiatkowski et al. (2019), closed-domain QA Clark et al. (2019), dialogue Cui et al. (2020), summarization Gliwa et al. (2019), named entity recognition Sang and Meulder (2003), and sentiment analysis Socher et al. (2013) are employed to understand the impact of model editing on the general abilities of LLMs.
Experimental results show that existing LLMs are not robust to weight perturbations, and editing even a few parameters can significantly affect their general abilities. Strikingly, with a single pass of editing involving less than 1% parameters, LLaMA-1 (7B) exhibited a drastic performance degradation to nearly 0 on all the tasks we tried. These results demonstrate that current editing algorithms struggle to work effectively in tandem with LLMs to simultaneously improve model factuality and maintain general abilities.
@@ 2 Related Work
Many studies have investigated model editing, including memory-based, meta-learning, and locate-then-edit Wang et al. (2024a); Yao et al. (2023). Memory-based methods do not modify model weights but store the editing facts with an external memory Mitchell et al. (2022b); Zhong et al. (2023). Mitchell et al. (2022b) stored edits in a base model and learned to reason over them to adjust its predictions as needed. The latter two classes of methods are developed to directly modify the internal parameters of models, which is the focus of this paper. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update model parameters Cao et al. (2021); Mitchell et al. (2022a). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022a) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly Dai et al. (2022); Meng et al. (2022, 2023). In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module. Besides, some works investigate the evaluation paradigm for model editing Zhong et al. (2023); Cohen et al. (2024); Ma et al. (2023); Li et al. (2024); Hase et al. (2023); Wu et al. (2023a); Gandikota et al. (2023); Wang et al. (2024b). For example, Cohen et al. (2024) introduced the ripple effects of editing, suggesting that editing a particular fact implies that many other facts need to be updated. Additionally, recent works have also applied editing in various domains, such as changing model personality Mao et al. (2023), editing multimodal models Cheng et al. (2023), protecting users privacy Wu et al. (2023b), etc.
@@
A main difference between this work and previous related studies should be highlighted. These approaches target at designing editing algorithms to improve or evaluation paradigms to assess the editing performance. In contrast, this study rethinks model editing and explores if current editing methods inadvertently cause the potential side effects on the underlying general abilities of LLMs. The contemporaneous work Gupta et al. (2024) presents a similar finding that model editing at scale leads to catastrophic forgetting, but no mitigation method is proposed. To the best of our knowledge, this paper makes the first call for attention to side effects on a variety of tasks beyond editing performance by presenting a systematical evaluation of four editing methods on three LLMs covering eight tasks. Besides, we also analyze the causes of side effects, and propose a regularization method to prevent editing overfitting.
@@ 3 Preliminary
To evaluate the effectiveness of editing methods, previous works focus on evaluation along three dimensions Cao et al. (2021); Mitchell et al. (2022a); Meng et al. (2022, 2023). First and foremost is reliability, aiming to ascertain the ability of the edited model to accurately recall the specific editing facts. The second dimension generalization seeks to validate the adaptability of the edited model by assessing its ability to recall the editing facts under diverse paraphrase prompts. The last dimension locality (a.k.a., specificity) is employed to verify the stability of the edited model by examining whether its output for unrelated inputs remains consistent after editing. Due to limited space, readers can refer to Appendix A for more detailed explanations and examples of the evaluation metrics.
@@ 4.1 Evaluation Paradigm
Single-editing involves examining the reliability and impact of making a single editing operation to a model. Specifically, it focuses on understanding how a model adapts to such a single alteration, and the implicit effect of such specific modifications on the overall performance. It is worth noting that a single editing operation can contain either only one editing instance or multiple ones in a batch, which is further discussed later. In practice, there are often situations where only a particular change is needed, so it’s crucial to understand how effectively the model integrates and preserves that individual edit. Therefore, evaluating the robustness to a single edit is crucial in determining its ability to retain the intended changes and overall performance.
In contrast to single-editing, multiple editing operations are conducted successively in sequential-editing Huang et al. (2023). Similarly, each editing operation in sequential-editing can also contain either only one editing instance or multiple ones in a batch. Ideally, models should retain the changes from previous edits when carrying out a new one Yao et al. (2023), which is decisive for the continual learning of future LLMs. Therefore, whether edited models can still maintain its general abilities after sequential editing is one of the important characteristics that should be considered. For this analysis, how the performance of edited models on a variety of tasks changes as the number of edits increases will be explored.
@@
The real world is ever-changing, so there is a huge amount of knowledge that needs to be dynamically added and updated into LLMs. Despite the effectiveness of many instance-editing methods Dai et al. (2022); Meng et al. (2022); Ma et al. (2023), ultimately at most a few dozens of pieces of knowledge can be updated Mitchell et al. (2022b), due to their relatively low but still non-negligible editing cost for a single instance. Since naive sequential applications of current state-of-the-art model editing methods fail to scale up Meng et al. (2023), one may wish to update hundreds or thousands of facts simultaneously in batch-editing. Notably, batch-editing can also be coupled with both the single- or sequential-editing settings.
@@
Zero-shot learning aims to solve tasks without labeled training examples, and recent studies have demonstrated the superiority of LLMs for zero-shot learning Brown et al. (2020); Wei et al. (2022); Chowdhery et al. (2023). Following these studies, we explore the zero-shot performance of unedited and edited models on a variety of tasks. Given a task instruction and a test problem that are concatenated as the input, the model is expected to generate a target text to address this problem. The instructions and input formats of different tasks are shown in Appendix B, which are taken from or inspired by Qin et al. (2023).
@@ 4.2 Evaluation Setup
Four popular editing methods were selected: (1) KN Dai et al. (2022) involved identifying neurons linked to knowledge expression using gradient-based attributions and then enhancing the MLP layer by adding scaled embedding vectors to those specific neurons. (2) MEND Mitchell et al. (2022a) learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form. (3) ROME Meng et al. (2022) involved localizing factual knowledge in a specific Transformer MLP layer and then updating it by directly writing new key-value pairs into the MLP module. (4) MEMIT Meng et al. (2023) expanded the capabilities of ROME by enabling the editing of large amounts of factual data through the updating of a sequence of MLP layers. It is notable that only MEND and MEMIT support batch-editing. All experiments were conducted using the EasyEdit tool Wang et al. (2024a), ensuring standardized and reproducible evaluation. All editing instances were randomly sampled from the editing dataset. Readers can refer to Appendix C for details of these editing methods.
@@
The popular model editing dataset Zero-Shot Relation Extraction (ZsRE) Levy et al. (2017) used in previous work Cao et al. (2021); Meng et al. (2022); Yao et al. (2023) was adopted in our experiments. ZsRE is a QA dataset using question rephrasings generated by back-translation as the equivalence neighborhood. Each input is a question about an entity, and plausible alternative edit labels are sampled from the top-ranked predictions of a BART-base model trained on ZsRE.
@@
Experiments were conducted on three LLMs: GPT-2 XL (1.5B) Radford et al. (2019), LLaMA-1 (7B) Touvron et al. (2023a), and LLaMA-2 (7B) Touvron et al. (2023b).
@@
To extensively explore whether model editing has side effects on the general abilities of LLMs, eight representative tasks were adopted: (1) Reasoning on the GSM8K Cobbe et al. (2021), and the results were measured by solve rate. (2) Natural language inference (NLI) on the RTE Dagan et al. (2005), and the results were measured by accuracy of two-way classification. (3) Open-domain QA on the Natural Question Kwiatkowski et al. (2019), and the results were measured by exact match (EM) with the reference answer after minor normalization as in Chen et al. (2017) and Lee et al. (2019). (4) Closed-domain QA on the BoolQ Clark et al. (2019), and the results were also measured by EM. (5) Dialogue on the MuTual Cui et al. (2020), and the results were measured by selecting one best-matched response from four available candidates, denoted as Recall4@1 as in Lowe et al. (2015). (6) Summarization on the SAMSum Gliwa et al. (2019), and the results were measured by the average of ROUGE-1, ROUGE-2 and ROUGE-L as in Lin (2004). (7) Named entity recognition (NER) on the CoNLL03 Sang and Meulder (2003), and the results were measured by entity-level F1-score. (8) Sentiment analysis on the SST2 Socher et al. (2013), and the results were measured by accuracy of two-way classification.